OpenStudy (anonymous):

write an expression for the number of ways , out of a standard 52-card deck , that 5-card hands can have 2 jacks and 3 queens.

OpenStudy (dumbcow):

(52C5)/(4C2)(4C3)(47C44) i think

OpenStudy (anonymous):

thank dumb cow

OpenStudy (anonymous):

can you solve for answer

OpenStudy (dumbcow):

yeah...you can use the combination feature on any scientific or graphing calculator as well

OpenStudy (anonymous):

ok,thank

OpenStudy (dumbcow):

i get 6.67 so i may be wrong, the number of ways should come out to a whole number

OpenStudy (dumbcow):

haha sorry i was over thinking the question: answer is just (4C2)*(4C3) = 6*4 = 24